# Quadratic Chabauty Experiments on Genus 2 Bielliptic Modular Curves in the LMFDB

### v1

@@ 4.1. Appearance of Weierstrass Points in the Rank 1 Case
As an example, consider the curve 60.48.2.c.1 and model $y^{2}=x^{6}+20x^{4}+25x^{2}-750$. This factors as $-2(x^{2}-5)(x^{2}+10)(x^{2}+15)$. The Weierstrass points live over $\mathbb{Q}(\sqrt{5}),\mathbb{Q}(\sqrt{-10})$, and $\mathbb{Q}\sqrt{-15})$. The primes used were 11, 13, 19, 23, and 37. As expected, we saw the points over $\mathbb{Q}(\sqrt{5})$ when using $p=11$ and 19, the points over $\mathbb{Q}(\sqrt{-10})$ using all five primes, and the points over $\mathbb{Q}(\sqrt{-15})$ using $p=19$ and 23.
@@ 5. Results
Any such curve has level divisible by that of the curve used in the main results. For example, our first curve is 14.48.2.g.1. We used the model $y^{2}=-x^{6}+83x^{4}-19x^{2}+1$. This model can also represent several other curves, including 28.96.2-14.g.1.1, 42.96.2-14.g.1.1, 56.96.2-14.g.1.1, and 70.96.2-14.g.1.1. We can add these as evidence for Conjectures 1.4 and 1.5.
@@ 6.1. $X_{ns}^{+}(15)$
We highlight points found on the modular curve $X_{ns}^{+}(15)$, which has LMFDB label 15.60.2.d.1. The model we use is $y^{2}=-45x^{6}+75x^{4}-15x^{2}+1$. The primes less than 100 of good, ordinary reduction for the curve are 7, 13, 19, 31, 37, 43, 61, 67, 73, and 97. For most of these primes, we recover no notable mock rational points. However, for $p\in\{19,31,61\}$ we observe four points: {(±$\{(\pm\frac{\sqrt{5}}{5},\pm\frac{4}{5})\}$. The sizes of the relevant outputs are in Table 2.
@@ 6.2. A Rank 1 Example
We highlight points over different number fields found on the modular curve 16.24.2.f.1. The model we use is $y^{2}=4x^{6}+6x^{4}+4x^{2}+1$. The Jacobian of this curve has rank 1. The results for two iterations of the analysis are in Table 3. Both cases recovered the two known rational points. Note that the points (±$(\pm i,\pm i)$ appeared both times, but the points over $\mathbb{Q}(\sqrt{-7})$ only appeared in the latter. A given viable field and prime will not necessarily capture all algebraic points.
@@
5 $\mathbb{Q}(\sqrt{-6})$ 66 (±$(\pm i,\pm i)$
37 $\mathbb{Q}(\sqrt{-7})$ 566 (±$(\pm i,\pm i),(\pm\frac{2\sqrt{-7}}{7},\pm\frac{5\sqrt{-7}}{49})$
@@ 6.3. A Return to $X_{0}(37)$
Using the model $y^{2}=-x^{6}-9x^{4}-11x^{2}+37$, we can run the analysis with $p=11$ over $\mathbb{Q}(\sqrt{-7})$. In this case, we observe the following set of points: {(±$\{(\pm\sqrt{-7},\pm 4),(\pm\frac{\sqrt{-7}}{3},\pm\frac{172}{27})\}$. The interested reader can find more about the quadratic points on $X_{0}(37)$ in Balakrishnan and Mazur’s 2024 paper regarding Ogg’s Torsion Conjecture [BM25] and an earlier paper of Box [Box21].
@@ 7. Future Work
First, a potential counterexample to the conjectures could be found in the following setup. Given a rational Heegner point on one of the elliptic curves $E_{1}$ or $E_{2}$ in the decomposition of the Jacobian of the genus 2 curve, this could pull back to a quadratic point on $X$. If this point satisfies the local height conditions required to be a mock rational point, it could be possible to appear in the quadratic Chabauty analysis while not having the properties of the points in the conjecture. These examples, if they exist, would be rare. It is worth exploring if they are possible, and to find them if so.
@@ 8. Appendix: Tables
Label Model Prime Number Field Points
14.48.2.g.1 $y^{2}=-x^{6}+83x^{4}-19x^{2}+1$ 37 $\mathbb{Q}(\sqrt{-3},\sqrt{21})$ (±$(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{21}}{9})$
15.60.2.d.1 $y^{2}=-45x^{6}+75x^{4}-15x^{2}+1$ 19 $\mathbb{Q}(\sqrt{5})$ $(\frac{\sqrt{5}}{5},\pm\frac{4}{5})$
16.48.2.a.1 $y^{2}=x^{6}-5x^{4}-5x^{2}+1$ 11 $\mathbb{Q}(\sqrt{-2})$ (±$(\pm 1,\pm 2\sqrt{2})$
18.72.2.f.1 $y^{2}=9x^{6}-99x^{4}+27x^{2}-1$ 7 $\mathbb{Q}(\sqrt{-3})$ (±$(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{-3}}{3})$
28.48.2.j.1 $y^{2}=x^{6}-83x^{4}+19x^{2}-1$ 19 $\mathbb{Q}(\sqrt{-3},\sqrt{7})$ (±$(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{-21}}{9})$
36.72.2.d.1 $y^{2}=-9x^{6}+99x^{4}-27x^{2}+1$ 13 $\mathbb{Q}(\sqrt{-3},\sqrt{3})$ (±$(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{3}}{3})$
48.48.2.dk.1 $y^{2}=12x^{6}+54x^{4}+48x^{2}+12$ 73 $\mathbb{Q}(\sqrt{3})$ $(0,\pm 2\sqrt{3})$
48.48.2.dm.1 $y^{2}=-24x^{6}+54x^{4}-24x^{2}+3$ 73 $\mathbb{Q}(\sqrt{3})$ $(0,\pm\sqrt{3})$
48.48.2.dw.1 $y^{2}=27x^{6}-72x^{4}+54x^{2}-8$ 73 $\mathbb{Q}(\sqrt{6})$ (±$(\pm\frac{\sqrt{6}}{9},\pm 2)$

### v2

@@ 4.1. Appearance of Weierstrass Points in the Rank 1 Case
Consider the curve 60.48.2.c.1 with model $y^{2}=x^{6}+20x^{4}+25x^{2}-750$ as an example. This factors as $2(x^{2}-5)(x^{2}+10)(x^{2}+15)$. The Weierstrass points live over $\mathbb{Q}(\sqrt{5}),\mathbb{Q}(\sqrt{-10})$, and $\mathbb{Q}\sqrt{-15})$. The primes used were 11, 13, 19, 23, and 37. As expected, we saw the points over $\mathbb{Q}(\sqrt{5})$ when using $p=11$ and 19, the points over $\mathbb{Q}(\sqrt{-10})$ using all five primes, and the points over $\mathbb{Q}(\sqrt{-15})$ using $p=19$ and 23.
@@ 5. Results
Any such curve has level divisible by that of the curve used in the main results. For example, our first curve is 14.48.2.g.1. In our analysis, we used the model $y^{2}=-x^{6}+83x^{4}-19x^{2}+1$. This model can also represent several other curves, including 28.96.2-14.g.1.1, 42.96.2-14.g.1.1, 56.96.2-14.g.1.1, and 70.96.2-14.g.1.1. We can add these as evidence for Conjecture 1.4.
@@ 6.1. $X_{ns}^{+}(15)$
We highlight points found on the modular curve $X_{ns}^{+}(15)$, which has LMFDB label 15.60.2.d.1. The model we use is $y^{2}=-45x^{6}+75x^{4}-15x^{2}+1$. The primes less than 100 of good, ordinary reduction for the curve are 7, 13, 19, 31, 37, 43, 61, 67, 73, and 97. For most of these primes, we recover no notable mock rational points. However, for $p\in\{19,31,61\}$ we observe four points: $\{(\pm\frac{\sqrt{5}}{5},\pm\frac{4}{5})\}$. The sizes of the relevant outputs are in Table 2.
@@ 6.2. A Rank 1 Example
We highlight points over different number fields found on the modular curve 16.24.2.f.1, using model $y^{2}=4x^{6}+6x^{4}+4x^{2}+1$. The Jacobian of this curve has rank 1. The results for two iterations of the analysis are in Table 3. Both cases recovered the two known rational points. Note that the points $(\pm i,\pm i)$ appeared both times, but the points over $\mathbb{Q}(\sqrt{-7})$ only appeared in the latter. A given viable field and prime will not necessarily capture all algebraic points.
@@
5 $\mathbb{Q}(\sqrt{-6})$ 66 $(\pm i,\pm i)$
37 $\mathbb{Q}(\sqrt{-7})$ 566 $(\pm i,\pm i),(\pm\frac{2\sqrt{-7}}{7},\pm\frac{5\sqrt{-7}}{49})$
@@ 6.3. A Return to $X_{0}(37)$
Using the model $y^{2}=-x^{6}-9x^{4}-11x^{2}+37$, we can run the analysis with $p=11$ over $\mathbb{Q}(\sqrt{-7})$. In this case, we observe the following set of points:
$\{(\pm\sqrt{-7},\pm 4),(\pm\frac{\sqrt{-7}}{3},\pm\frac{172}{27})\}.$
The interested reader can find more about the quadratic points on $X_{0}(37)$ in Balakrishnan and Mazur’s 2024 paper regarding Ogg’s Torsion Conjecture [BM25] and an earlier paper of Box [Box21].
@@ 7. Future Work
First, a potential counterexample to the conjecture could be found in the following setup. Given a rational Heegner point on one of the elliptic curves $E_{1}$ or $E_{2}$ in the decomposition of the Jacobian of the genus 2 curve, this could pull back to a quadratic point on $X$. If this point satisfies the local height conditions required to be a mock rational point, it could be possible to appear in the quadratic Chabauty analysis while not having the properties of the points in the conjecture. These examples, if they exist, would be rare. It is worth exploring if they are possible, and to find them if so.
@@ 8. Appendix: Tables
Label Model $y^{2}=f(x)$ $p$ Number Field Points
14.48.2.g.1 $x^{6}+83x^{4}-19x^{2}+1$ 37 $\mathbb{Q}(\sqrt{-3},\sqrt{21})$ $(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{21}}{9})$
15.60.2.d.1 $45x^{6}+75x^{4}-15x^{2}+1$ 19 $\mathbb{Q}(\sqrt{5})$ $(\frac{\sqrt{5}}{5},\pm\frac{4}{5})$
16.48.2.a.1 $x^{6}-5x^{4}-5x^{2}+1$ 11 $\mathbb{Q}(\sqrt{-2})$ $(\pm 1,\pm 2\sqrt{2})$
18.72.2.f.1 $9x^{6}-99x^{4}+27x^{2}-1$ 7 $\mathbb{Q}(\sqrt{-3})$ $(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{-3}}{3})$
28.48.2.j.1 $x^{6}-83x^{4}+19x^{2}-1$ 19 $\mathbb{Q}(\sqrt{-3},\sqrt{7})$ $(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{-21}}{9})$
36.72.2.d.1 $9x^{6}+99x^{4}-27x^{2}+1$ 13 $\mathbb{Q}(\sqrt{-3},\sqrt{3})$ $(\pm\frac{\sqrt{-3}}{3},\pm\frac{8\sqrt{3}}{3})$
48.48.2.dk.1 $12x^{6}+54x^{4}+48x^{2}+12$ 73 $\mathbb{Q}(\sqrt{3})$ $(0,\pm 2\sqrt{3})$
48.48.2.dm.1 $24x^{6}+54x^{4}-24x^{2}+3$ 73 $\mathbb{Q}(\sqrt{3})$ $(0,\pm\sqrt{3})$
48.48.2.dw.1 $27x^{6}-72x^{4}+54x^{2}-8$ 73 $\mathbb{Q}(\sqrt{6})$ $(\pm\frac{\sqrt{6}}{9},\pm 2)$